OpenStudy (anonymous):

Determine the sum of all the multiples of 4 between 1 and 999. Could you please tell me how to find n from here: n = d = 4? t1 = 4? Sn = n/2 [2 (t1) + (n-1) d]

hartnn (hartnn):

d= 4, t1 = 4 are correct, can you tell me what the last term would be ?? last number less than 999, that is divisible by 4 ?

OpenStudy (anonymous):

996

OpenStudy (anonymous):

is that okay?

hartnn (hartnn):

yes, thats correct :) so, 4 is the 1st term, 8 is the 2nd term, so in general '4n' would be the n'th term, right ? (n=1, t1=4, n=2, t2=8 and so on) got this ? so, your "n" would be found from 4n = 996, n=... ?

OpenStudy (anonymous):

249

OpenStudy (anonymous):

why is 4n the nth term?

hartnn (hartnn):

see, the terms are 4,8,12,16,20,......996 right ? so if you observe that those are multiples of 4 ! like 1st term = 4 *1 = 4 2nd term = 4*2 = 8 3rd term = 4*3 = 12 .... nth term = 4*n =4n

OpenStudy (anonymous):

Thank you. I get it now.

hartnn (hartnn):

welcome ^_^